OpenStudy (anonymous):

Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature.

OpenStudy (dangerousjesse):

It's right :)

OpenStudy (squirrels):

So you want us to write your essay?

OpenStudy (anonymous):

??? @dan815 @SolomonZelman help I am doing a essay

OpenStudy (solomonzelman):

Do you know the formula for converting Fº to Cº ?

OpenStudy (anonymous):

that up there is the question and I need to answer it but I don't know how whats the answer??

OpenStudy (johnweldon1993):

F to C would be \[\large \frac{5}{9}(F - 32)\] So for example..if Fahrenheit would be 32 degrees... \[\large \frac{5}{9}(32 - 32) = 0\] and that would be 0 degrees Celsius

OpenStudy (anonymous):

thank u ;-)

OpenStudy (johnweldon1993):

No problem! :)